OpenStudy (anonymous):

Fifteen people volunteer for a park cleanup. The number of volunteers increases by 7 people each month for several months. After how many months will there be 50 volunteers?

OpenStudy (anonymous):

Write an equation: 15+7x=50, where x is the number of months. Then solve for x

OpenStudy (anonymous):

I do hope you know how I got to the equation...

OpenStudy (anonymous):

don't you like put negative 50 at the bottom of 15 and 50? or?

OpenStudy (anonymous):

Setting up the equation: You start with 15 people, so those are constant. Since the people are increasing by 7 every month, you would add 7*# of months to 15 to get the total amount of people. Solving the equation: You have a target, 50 people. 15+7x=50 You want to isolate the x variable, so you can solve in two steps. 1) subtract 15 from both sides: 7x=35 2)divide by 7: x=35/7 = 5 Therefore, after 5 months, the number of people will reach 50 people. Sorry for the long post.

OpenStudy (anonymous):

alright that made more sense. thank you for the help.